Operations Report – December 20th

Crew 185
Operations Report Wednesday 12/20/2017
SOL: 2
Name of person filing report: A.Passeron
Non-nominal systems:
– Spacesuit 3 : The ventilation system of spacesuit N°3 (mine) stopped during the engineering EVA. Back to the EVA room at 11:36am, I plugged the backpack but the DEL of the charger was still turned off. I meticulously moved the plug of the charger and the red DEL of the charger switched on.  At 2:30pm, I turned the ventilation system on and it was working properly. There is thus a faulty connection. I will open the backpack this evening to see whether or not a wire is badly connected to the outlet of the spacesuit.
– Spacesuit 4 : Back from the EVA at 1:25pm, while undressing her spacesuit, Ilaria pulled out the right ventilation tube of her backpack. I just had to plug it back into the spacesuit.
– Hot water circuit of the sink : There is not hot water circulating in the sink.
Generator (hours run):
Generator turned off at 8:30 am.
Generator turned on at 5:15pm.
Solar— SOC @ 5:15pm : 78%
Diesel – 80%
Propane – 58 percent volume
Ethanol Free Gasoline (5 Gallon containers for ATV) – 3.5 Gallons
Water (trailer) – 50 gallons
Water (static) – 350 gallons
Trailer to Static Pump used – Yes
Static to Loft Pump used – Yes
Water Meter: 50 gallons
Toilet tank emptied: No
ATVs Used: No
Oil Added? No
ATV Fuel Used:
# Hours the ATVs were Used today: 00:00 hours
Notes on ATVs: ATVs were nominal.
Deimos rover used: No
Hours: 00:00 hours
Beginning charge:
Ending charge:
Currently charging:
Sojourner rover used: Assigned to director only.
Hours: 00:00 Hours
Beginning charge:
Ending charge:
Currently charging:
Spirit rover used:  No
Hours: 00:00 Hours
Beginning charge:
Ending charge:
Currently charging: Yes
Opportunity rover used:  No
Hours: 00:00 Hours
Beginning charge:
Ending charge:
Currently charging: Yes
Curiosity rover used:  No
Hours:
Beginning charge:
Ending charge:
Currently charging:
HabCar used and why, where?  No
Summary of internet: All Nominal
Summary of suits and radios: Spacesuit N°4 repaired. The outlet of spacesuit N°3 has a faulty connection.
Summary of Hab operations:  There is not hot water circulating in the sink
Summary of GreenHab operations: All nominal
Summary of ScienceDome operations: All nominal
Summary of RAMM operations: Not Operational
Summary of health and safety issues: Crew is Healthy
Questions, concerns and requests to Mission Support: